OpenStudy (anonymous):

Need help for trigo.. fan and medal given! Show that [2(cos x)^2 - 1] (sec x)^2 = 2 - (sec x)^2

OpenStudy (sidsiddhartha):

ok first u need to know this basic indentity \[secx=\frac{ 1 }{ cosx }\]

OpenStudy (anonymous):

\[(2\cos ^{2}x-1)\sec ^{2}x = 2-\sec ^{2}x\]

OpenStudy (sidsiddhartha):

are u familiar with this?

OpenStudy (anonymous):

i know that

OpenStudy (anonymous):

ya o.o

OpenStudy (sidsiddhartha):

ok so just multiply

OpenStudy (anonymous):

i did..

OpenStudy (anonymous):

=.=

OpenStudy (sidsiddhartha):

\[(2\cos^2x-1)\sec^2x=2\cos^2x.\sec^2x-\sec^2x\] right?

OpenStudy (anonymous):

oh ok i got it

OpenStudy (anonymous):

thanks ! :)

OpenStudy (sidsiddhartha):

yw!!!

OpenStudy (anonymous):

fanned and medaled ! :D thank yew!

OpenStudy (sidsiddhartha):

happy to help ya :)